# On the Similarities of Embeddings in Contrastive Learning

### v1

@@ 7 Conclusion
To understand contrastive learning (CL), we mathematically analyzed the distributions of similarities of embeddings, measured for positive pairs and negative pairs. Our theoretical results in full-batch CL demonstrate that misalignment of positive pairs becomes inevitable when the average similarity of negative pairs falls below its optimal value, a situation that can arise with existing contrastive loss functions. In mini-batch CL, we prove that the variance of negative-pair similarities increases as the batch size decreases. To address the detrimental variance introduced by mini-batching, we propose an auxiliary loss that reduces this variance, which also enhances the performance of existing CL methods.
We aim to extend this work in two directions. First, we seek to disentangle the variance of negative-pair similarities that reflects intrinsic data structure from that caused by mini-batching, enabling us to target only the variance that degrades representation quality. Second, we plan to analyze the behavior of embedding similarities not only during pretraining but also during fine-tuning, to gain deeper insights into its dynamics throughout different stages of training.
@@ Acknowledgements
Appendix A Loss Functions
We outline how various loss functions commonly used in CL can be instantiated by the general formulation provided in Def. 3.1 and Def. 3.2. For each case, we specify the corresponding choices of functions and parameters.
@@ Appendix A Loss Functions
Figure 5: Illustration comparing four different contrastive loss functions, all following the form of Def. 3.1. The green area represents the positive pair, while the blue-striped regions indicate the negative pairs that are normalized together with the positive pair for each loss.
@@ C.3 Proofs for Full-Batch CL
Suppose that $d\geq n-1$. Let the contrastive loss function $\mathcal{L}\left({\bm{U}},{\bm{V}}\right)$ be one of the following forms.
@@
Suppose that $d\geq n$. Let the contrastive loss function $\mathcal{L}\left({\bm{U}},{\bm{V}}\right)$ be the form of $\mathcal{L}_{\operatorname{ind-add}}\left({\bm{U}},{\bm{V}}\right)$ in Def. 3.2, where $(c_{1},c_{2})=(1,0)$ and $\phi^{\prime}\left(1\right)<\frac{n-2}{2(n-1)}\cdot\psi^{\prime}\left(-\frac{1}{n-1}\right)$. Then, embedding similarities for the full-batch optimal encoder $f^{\star}$ in (1) satisfy
@@
We prove for each category of loss function, $\mathcal{L}_{\operatorname{info-sym}}\left({\bm{U}},{\bm{V}}\right)$ in Def. 3.1 and $\mathcal{L}_{\operatorname{ind-add}}\left({\bm{U}},{\bm{V}}\right)$ in Def. 3.2, separately.
@@
embedding similarities of the (full-batch) optimal encoder $f^{\star}$ in (1) satisfy
@@
following from Theorem C.10, the similarities of the (full-batch) optimal encoder $f^{\star}$ in (1) satisfy
@@ C.4 Proofs for Mini-Batch CL
Strictly speaking, the notation ${\bm{U}}\cup{\bm{V}}$ should be interpreted as a concatenation of arrays rather than a set union, as the ordering of vectors is relevant in the computation.
@@
Suppose $d\geq m-1$. Let the contrastive loss function $\mathcal{L}\left({\bm{U}},{\bm{V}}\right)$ is one of the forms in Theorem 5.1. Define $f^{\star}_{\operatorname{batch}}$ as the optimal encoder that minimizes the fixed mini-batch loss, given by
@@ Appendix E Additional Experiments on the ImageNet Dataset
We further validate the effectiveness of the proposed loss term through experiments on the ImageNet dataset. Table 4 reports the top-1 and top-5 classification accuracies on ImageNet-100 for various temperature values. For each temperature, we compare SimCLR with and without the proposed loss. The results demonstrate that incorporating our loss generally improves performance across a range of temperatures, with the largest gain observed at $t=0.3$.
Table 4: Effect of our proposed loss term (when combined with SimCLR) on the top-1 and top-5 classification accuracies (%). Bold entries indicate the highest accuracy.
@@ Appendix F Discussion on the Proposed Loss for Variance Reduction
The auxiliary loss term $\mathcal{L}_{\operatorname{VRNS}}(\mathbf{U},\mathbf{V})$ proposed in Def. 5.7 shows effectiveness in the following scenarios:

### v2

@@ 7 Conclusion
To understand contrastive learning (CL), we mathematically analyze the distributions of similarities of embeddings, measured for positive pairs and negative pairs. Our theoretical results in full-batch settings demonstrate that misalignment of positive pairs becomes inevitable when the average similarity of negative pairs falls below its optimal value, a situation that can arise with existing contrastive losses. In mini-batch settings, we prove that the variance of negative-pair similarities increases as the batch size decreases—a distinctive characteristic absent in full-batch settings and a potential contributor to the performance degradation observed in mini-batch settings. To address this, we propose an auxiliary loss that explicitly reduces the variance of negative-pair similarities. Empirical results show that incorporating the proposed loss improves the performance of CL methods, especially in small-batch settings.
Promising directions for future work include extending this work in two directions. First, disentangling the variance of negative-pair similarities that reflects intrinsic data structure from that caused by mini-batching could enable targeting only the variance that degrades representation quality. Second, analyzing the behavior of embedding similarities not only during pretraining but also during fine-tuning may provide deeper insights into its dynamics throughout different stages of training.
@@ Acknowledgements
Appendix A Contrastive Losses
We outline how various losses commonly used in CL can be instantiated by the general formulation provided in Def. 3.1 and Def. 3.2. For each case, we specify the corresponding choices of functions and parameters.
@@ Appendix A Contrastive Losses
Figure 5: Illustration comparing four different contrastive losses, all following the form of Def. 3.1. The green area represents the positive pair, while the blue-striped regions indicate the negative pairs that are normalized together with the positive pair for each loss.
@@ C.3 Proofs for Full-Batch CL
Suppose that $d\geq n-1$. Let the contrastive loss $\mathcal{L}\left({\bm{U}},{\bm{V}}\right)$ be one of the following forms.
@@
Suppose that $d\geq n$. Let the contrastive loss $\mathcal{L}\left({\bm{U}},{\bm{V}}\right)$ be the form of $\mathcal{L}_{\operatorname{ind-add}}\left({\bm{U}},{\bm{V}}\right)$ in Def. 3.2, where $(c_{1},c_{2})=(1,0)$ and $\phi^{\prime}\left(1\right)<\frac{n-2}{2(n-1)}\cdot\psi^{\prime}\left(-\frac{1}{n-1}\right)$. Then, embedding similarities for the full-batch optimal encoder $f^{\star}$ in (1) satisfy
@@
We prove for each category of loss, $\mathcal{L}_{\operatorname{info-sym}}\left({\bm{U}},{\bm{V}}\right)$ in Def. 3.1 and $\mathcal{L}_{\operatorname{ind-add}}\left({\bm{U}},{\bm{V}}\right)$ in Def. 3.2, separately.
@@
embedding similarities of the full-batch optimal encoder $f^{\star}$ in (1) satisfy
@@
following from Theorem C.10, the similarities of the full-batch optimal encoder $f^{\star}$ in (1) satisfy
@@ C.4 Proofs for Mini-Batch CL
Here, ${\bm{U}}\cup{\bm{V}}$ denotes the concatenation of the two sets, representing the full collection of all $n+m$ vectors from ${\bm{U}}$ and ${\bm{V}}$.
@@
Suppose $d\geq m-1$. Let the contrastive loss $\mathcal{L}\left({\bm{U}},{\bm{V}}\right)$ be one of the forms in Theorem 5.1. Define $f^{\star}_{\operatorname{batch}}$ as the optimal encoder that minimizes the fixed mini-batch loss, given by
@@ Appendix E Additional Experiments on the ImageNet Dataset
Table 4: Effect of our proposed loss (when combined with SimCLR) on the top-1 and top-5 classification accuracies (%). Bold entries indicate the highest accuracy.
@@
We further validate the effectiveness of the proposed loss through experiments on the ImageNet dataset. Table 4 reports the top-1 and top-5 classification accuracies on ImageNet-100 for various temperature values. For each temperature, we compare SimCLR with and without the proposed loss. The results demonstrate that incorporating our loss generally improves performance across a range of temperatures, with the largest gain observed at $t=0.3$.
@@ Appendix F Discussion on the Proposed Loss for Variance Reduction
The auxiliary loss $\mathcal{L}_{\operatorname{VRNS}}(\mathbf{U},\mathbf{V})$ proposed in Def. 5.7 shows effectiveness in the following scenarios: